OpenStudy (anonymous):

Find dy/dx if y = ln[(square root of x)/(5 - x)]

OpenStudy (anonymous):

\[y=\frac{\ln \sqrt{x}}{5-x}=\frac{\frac{1}{2}\ln x}{5-x}=\frac{\ln x}{10-2x}\] use this rule \[(\frac{f}{g})'=\frac{f'g-fg'}{g^2}\]

OpenStudy (anonymous):

@Jonask, it looks like the log contains all of \(\dfrac{\sqrt x}{5-x}\).

OpenStudy (anonymous):

I would make use of log properties: \[y=\ln\left[\frac{\sqrt x}{5-x}\right]\\ ~~=\frac{1}{2}\ln x-\ln(5-x)\]

OpenStudy (anonymous):

ohhh sry

OpenStudy (anonymous):

but your right thats the way we shud do it

OpenStudy (anonymous):

Quotient rule can get messy pretty quickly :)

OpenStudy (anonymous):

now use \[\frac{ d }{ dx }\ln f(x)=\frac{ f'(x) }{ f(x) }\]

OpenStudy (anonymous):

from here \[\frac{ 1 }{ 2 }\ln x-\ln (5-x)\]

OpenStudy (anonymous):

The second one would be 1/(5-x), right? I'm just not sure what to do with the first one because of its 1/2. Would it just be 1/2 * (1/x)? Or am I doing it entirely wrong? xD

OpenStudy (anonymous):

yes your right you do it as if the 1/2 is not there 1/2*(1/x) is correct

OpenStudy (anonymous):

Well, I can't find anyway of simplifying it without making a huge mess, so do ya think y = (1/2x) - (1/5 - x) would be an acceptable final answer?

OpenStudy (anonymous):

yes \[\frac{ 1 }{ 2x }-\frac{ 1 }{ 5-x }=\frac{ 5+x }{ 10x-2x^2}\]

OpenStudy (anonymous):

Alright. Thank you very much! :D

OpenStudy (anonymous):

i think ur missing a "-" sign there derivative of ln(5-x)=-1/(5-x)

OpenStudy (anonymous):

yes